Wavelength of a light beam is 25 micrometre , it`s wave number is-(1)4.0×10^2 cm^-1 (2)4.0×10^6 cm^-1 (3)4.0×10^6 m^-1 (4)25.0×10^4 cm^-1Explain???

Grade:11

1 Answers

Kartik
7 years ago
25 micrometer =25*10^-6 m=25*10-4 cmWavenumber is receprokal of wavelengthWave number=1/25*10^-4=0.04*10^-4=4*10^2Hence options 1 is correct